# ChatGPT’s One-year Anniversary: Are Open-Source Large Language Models Catching up?

### 28 Nov 2023 v1

@@ Abstract
**footnotetext: Authors contributed equally and are ranked by alphabetical order.
@@ 1 Introduction
However, since ChatGPT is not open-sourced and its access is controlled by a private company, most of its technical details remain unknown. Despite the claim that ChatGPT follows the procedure introduced in InstructGPT (also called GPT-3.5) (Ouyang et al., 2022b), its exact architecture, pre-training data and fine-tuning data are unknown. Such close-source nature generates several key issues. First, without knowing the internal details such as the pretraining and finetuning procedure, it is hard to properly estimate its potential risks to the society, especially knowing that LLMs can notoriously generate toxic, unethical and untruthful content. Second, it has been reported that ChatGPT’s performance changes over time hindering reproducible results (Chen et al., 2023c). Third, ChatGPT has experienced multiple outages, with two major ones only in November 2023 during which the access to ChatGPT website and its API was completely blocked. Finally, enterprises adopting ChatGPT may be concerned with the heavy cost of calling APIs, service outages, data ownership and privacy issues, and unpredictable events such as the recent boardroom drama about the CEO Sam Altman’s dismissal to staff rebellion, and his eventual return (REUTERS source).
(a) Open-source LLMs vs. ChatGPT on MT-bench Zheng et al. (2023), ELO Rating, MMLU (Hendrycks et al., 2021b) and AlpacaEval Li et al. (2023d).
@@
Open-source LLMs consist in a promising direction as they can potentially remediate or bypass most of the aforementioned issues. For this reason, the research community has been pushing for maintaining high-performing LLMs in open-source. As it stands today, it is however widely believed that open-source LLMs such as Llama-2 (Touvron et al., 2023b) or Falcon (Almazrouei et al., 2023) lag behind their closed-source counterparts such as OpenAI’s GPT3.5 (ChatGPT) and GPT-4 (OpenAI, 2023b), Anthropic’s Claude [^2] https://www.anthropic.com/index/introducing-claude or Google’s Bard [^3] https://blog.google/technology/ai/bard-google-ai-search-updates, with GPT-4 generally assumed to champion them all as of late 2023. However, the gap is getting narrower and narrower, and open-source LLMs are quickly catching up. In fact, as it is illustrated in Figure 1 on certain tasks, the best open-source LLMs already perform better than GPT-3.5-turbo. Yet, it is not a straightforward uphill battle for open-source LLMs. The LLMs landscape is constantly evolving: closed-source LLMs are re-trained on newer data regularly, open-source LLMs are released almost weekly, and there is a myriad of evaluation datasets and benchmarks being used to compare LLMs, making singling out a best LLM especially challenging.
In this survey, we aim to consolidate recent papers of open-source LLMs and provide an overview of open-source LLMs that match or surpass ChatGPT in various domains. Our contributions are three-fold:
@@
In the following article, we start by introducing background prerequisites (Section 2), then provide an in-depth review of open-source LLMs that beat ChatGPT in various domains (Section 3), followed by a discussion on insights and issues of open-source LLMs (Section 4), finally we conclude the survey (Section 5).
{forest} for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, text width=2.7cm, s sep=6pt, calign=edge midpoint, , [LLM’s Capabilities, fill=gray!45, parent [General §3.1, domain_box [AlpacaEval, MT-bench, ELO rating, Open-LLM leaderboard, datasets [Llama-2, WizardLM, GodziLLa2, Zephyr, models] ] ] [Agent §3.2, domain_box [Using Tools, domain_box [API-Bank, ToolBench, APIBench, ToolAlpaca, etc., datasets [Gorilla, ToolLLaMA, few_models] ] ] [Self-Debugging, domain_box [InterCode-Bash, InterCode-SQL, MINT-MBPP, MINT-HumanEval, RoboCodeGen, etc., datasets ] ] [Following NL Feedback, domain_box [MINT, few_datasets] ] [Exploring Environments, domain_box [ALFWorld, InterCode-CTF, WebArena, datasets] ] [Lemur-chat, AgentLlama, models] ] [Logical Reasoning §3.3, domain_box [Math, domain_box [GSM8K, MATH, TheoremQA, etc., datasets [WizardMath, few_models] ] ] [Coding, domain_box [HumanEval, MBPP, APPs, etc., datasets [WizardCoder, few_models] ] ] ] [Modelling Long-context §3.4, domain_box [SCROLLS, Zero-SCROLLS, LongBench, L-Eval, BAMBOO, M4LE, etc., datasets [Llama-2-long, few_models] ] ] [Application-specific §3.5.5, domain_box [Query-focused Summarization, domain_box [QMSum, SQuALITY, CovidET, NEWTS, datasets [finetuned model, few_models] ] ] [Open-ended QA, domain_box [NQ,TriviaQA,NewsQA,SQuAD,Quoref, NarrativeQA,DROP, datasets [InstructRetro, few_models] ] ] [Medical, domain_box [Dreaddit, loneliness, MIMIC-CXR, OpenI, etc., datasets [MentaLLaMA, Radiology-Llama-2, few_models] ] ] [Generating Structured Data, domain_box [Rotowire, Struc-Bench-Latex, Struc-Bench-HTML, datasets [Struct-Bench, few_models] ] ] [Generating Critiques, domain_box [AlpacaFarm, FairEval, CritiqueEval, etc., datasets [Shepherd, few_models] ] ] ] [Trustworthiness §3.6, domain_box [Hallucination, domain_box [TruthfulQA, FactualityPrompt, FActScore, KoLA-KC, HaluEval, FACTOR, datasets [Platypus, Chain-of-Verification, etc., few_models] ] ] [Safety, domain_box [SafetyBench, XSTEST, etc., datasets] ] ] ]
@@ 2.1 Training Regimes
All LLMs rely on large-scale self-supervised pre-training on Internet text data (Radford et al., 2018; Brown et al., 2020). Decoder-only LLMs follow the causal language modeling objective, through which the model learns to predict the next token conditioning on the sequence of previous tokens (Bengio et al., 2000). As per pre-training details shared by open-source LLMs (Touvron et al., 2023a), sources of text data include CommonCrawl [^5] https://commoncrawl.org, C4 (Raffel et al., 2020), GitHub, Wikipedia, books, and online discussion exchanges such as Reddit or StackOverFlow. It is widely acknowledged that scaling the size of pre-training corpus improves the model performance, and works hand-in-hand with scaling the model size, a phenomenon referred to as scaling laws, and analyzed in depth in (Hoffmann et al., 2022a). Modern-day LLMs pre-train on a corpus from hundreds of billions to several trillions of tokens (Touvron et al., 2023b; Penedo et al., 2023).
@@
Fine-tuning aims to adapt a pre-trained LLM to a downstream task, by updating weights with the available supervision, which usually forms a dataset orders of magnitude smaller than the one used for pre-training (Devlin et al., 2018). T5 (Raffel et al., 2020) was among the first to frame fine-tuning into a text-to-text unified framework, with natural language instructions describing each task. Instruction-tuning later extended fine-tuning by training jointly on several tasks (Wei et al., 2021a; Aribandi et al., 2021), each described with natural language instructions. Instruction-tuning quickly gained in popularity, due to its ability to drastically improve zero-shot performance of LLMs, including on new tasks (unseen during training), and especially at larger models scale. Standard instruction-tuning with multi-task supervised fine-tuning may still not result in models following humans intentions, and can be improved with Reinforcement Learning from Human Feedback (RLHF): human annotators rank outputs from the fine-tuned model, which are used to fine-tune again with reinforcement learning (Ouyang et al., 2022b). Recent work showed that humans feedback may be replaced with feedback from an LLM, a process referred to as Reinforcement Learning from AI Feedback (RLAIF) Bai et al. (2022b). A line of work focuses on quality over quantity when building an instruction-tuning dataset of diverse tasks: Lima (Zhou et al., 2023a) outperforms GPT-3 with a Llama-65B fine-tuned on just 1,000 examples, and Alpagasus (Chen et al., 2023b) improves on Alpaca (Taori et al., 2023) by cleaning its instruction fine-tuning dataset from 52k to 9k examples.
@@
Continual pre-training consists in performing another round of pre-training from a pre-trained LLM, typically with a lesser volume of data than in the first stage. Such process may be useful to elicit new properties in the LLM. For instance, continual pre-training is used in Lemur (Xu et al., 2023d) to improve coding and reasoning capacities, and Llama-2-long (Xiong et al., 2023) to extend context window.
@@
There exists several alternative methods for sequence generation with auto-regressive decoding with a LLM, which differ by the degree of randomness and diversity in the output. Increasing the temperature during sampling makes outputs more diverse, while setting it to 0 falls back to greedy decoding, which may be needed in scenarios necessitating deterministic outputs. Sampling methods top-k (Fan et al., 2018) and top-p (Holtzman et al., 2019) constrain the pool of tokens to sample from at each decoding step.
@@ 2.2 Task Domains and Evaluation
3 Open-Source LLMs vs ChatGPT
@@ 3.1 General Capabilities
As numerous LLMs are released week upon week, each claiming superior performance, it becomes challenging to identify true advancements and the leading model. Therefore, it is crucial to comprehensively assess the performance of these models across a broad spectrum of tasks to understand their general capabilities. This section covers benchmarks using LLM-based (e.g., GPT-4) and traditional (e.g., ROUGE (Lin, 2004) and BLEU (Papineni et al., 2002)) evaluation metrics.
  • MT-Bench (Zheng et al., 2023) is designed to test multi-turn conversation and instruction-following ability from eight perspectives, including writing, roleplay, extraction, reasoning, math, coding, knowledge I (STEM), and knoweldge II (humanities/social science). Stronger LLMs are utilized as judges to evaluate the models for this benchmark.
  • AlpacaEval (Li et al., 2023d) is an LLM-based automatic evaluator based on AlpacaFarm (Dubois et al., 2023) evaluation set, which tests the ability of models to follow general user instructions. It benchmarks candidate models against Davinci-003 responses utilizing stronger LLMs (e.g., GPT-4 and Claude), which generate the candidate model’s win rate.
@@
  • FairEval-Vicuna (Wang et al., 2023b) utilizes both multiple evidence calibration and balanced position calibration on a set of 80 questions from the Vicuna Benchmark (Zheng et al., 2023). FairEval-Vicuna offers a more impartial evaluation outcome within the paradigm of adopting LLMs as evaluators, which closely aligns with human judgements.
Models MT-Bench AlpacaEval Open LLM Leaderboard
Llama-2-70B-chat 6.86 92.66 -
WizardLM-70B 7.71 92.91 57.17
GodziLLa2-70B - - 67.01
Zephyr-7B 7.34 90.60 52.15
GPT-3.5-turbo 7.94 81.71 70.21
GPT-4 8.99 95.28 85.36
@@
Models
Llama-2-70B (Touvron et al., 2023b), a prominent open-source LLM, has been pre-trained on a massive dataset of two trillion tokens. It demonstrates remarkable results across various general benchmarks. When further fine-tuned with instruction data, the Llama-2-chat-70B variant exhibits enhanced capabilities in general conversational tasks. In particular, Llama-2-chat-70B achieves a 92.66% win rate in AlpacaEval, surpassing the performance of GPT-3.5-turbo by 10.95%. Nonetheless, GPT-4 remains the top performer among all LLMs with a win rate of 95.28%. Zephyr-7B (Tunstall et al., 2023), another smaller model, uses distilled direct preference optimization (Rafailov et al., 2023a) and achieves comparable results to 70B LLMs on AlpacaEval with a win rate of 90.6%. It even surpasses Llama-2-chat-70B on MT-Bench, scoring 7.34 against 6.86. Additionally, WizardLM-70B (Xu et al., 2023a) has been instruction fine-tuned using large amounts of instruction data with varying levels of complexity. It stands out as the highest-scoring open-sourced LLM on MT-Bench with a score of 7.71. However, this is still slightly lower than the scores of GPT-3.5-turbo (7.94) and GPT-4 (8.99). Although Zephyr-7B shows top performance in the MT-Bench, it falls short in the Open LLM Leaderboard, scoring only 52.15%. On the other hand, GodziLLa2-70B (Philippines, 2023), an experimental model that combines various proprietary LoRAs from Maya Philippines  [^6] Maya (https://www.maya.ph) is a Filipino financial services and digital payments company. and the Guanaco Llama 2 1K dataset (mlabonne, 2023) with Llama-2-70B, achieves a more competitive score of 67.01% on the Open LLM Leaderboard. This performance is comparable to that of GPT-3.5-turbo, which scores 70.21%. However, both are still notably behind GPT-4, which leads with a substantial score of 85.36%. UltraLlama (Ding et al., 2023) utilizes fine-tuning data with enhanced diversity and quality. It matches GPT-3.5-turbo’s performance in its proposed benchmark while exceeding it in areas of world and professional knowledge.

### 15 Jan 2024 v4

@@ Abstract
**footnotetext: Authors contributed equally and are ranked by alphabetical order.footnotetext: Work done full-time at Salesforce
@@ 1 Introduction
However, since ChatGPT is not open-sourced and its access is controlled by a private company, most of its technical details remain unknown. Despite the claim that it follows the procedure introduced in InstructGPT (also called GPT-3.5) (Ouyang et al., 2022b), its exact architecture, pre-training data and fine-tuning data are unknown. Such close-source nature generates several key issues. First, without knowing the internal details such as the pre-training and fine-tuning procedure, it is hard to properly estimate its potential risks to the society, especially knowing that LLMs can notoriously generate toxic, unethical and untruthful content. Second, it has been reported that ChatGPT’s performance changes over time hindering reproducible results (Chen et al., 2023d). Third, ChatGPT has experienced multiple outages, with two major ones only in November 2023 during which the access to ChatGPT website and its API was completely blocked. Finally, enterprises adopting ChatGPT may be concerned with the heavy cost of calling APIs, service outages, data ownership and privacy issues, and other unpredictable events such as the recent boardroom drama about the CEO Sam Altman’s dismissal to staff rebellion, and his eventual return (REUTERS source).
(a) Open-source LLMs vs. ChatGPT on MT-bench Zheng et al. (2023), ELO Rating, MMLU (Hendrycks et al., 2021b) and AlpacaEval Li et al. (2023e).
@@
Open-source LLMs, on the other hand, offer a promising direction as they can potentially remediate or bypass most of the aforementioned issues. For this reason, the research community has been actively pushing for maintaining high-performing LLMs in open-source. However, as it stands today (as of late 2023), it is widely believed that open-source LLMs such as Llama-2 (Touvron et al., 2023b) or Falcon (Almazrouei et al., 2023) lag behind their closed-source counterparts such as OpenAI’s GPT3.5 (ChatGPT) and GPT-4 (OpenAI, 2023b), Anthropic’s Claude [^2] https://www.anthropic.com/index/introducing-claude or Google’s Bard [^3] https://blog.google/technology/ai/bard-google-ai-search-updates, with GPT-4 generally assumed to champion them all. However, what is very encouraging is that the gap is getting narrower and narrower, and open-source LLMs are quickly catching up. In fact, as it is illustrated in Figure 1, the best open-source LLMs already perform better than GPT-3.5-turbo on some standard benchmarks. Yet, it is not a straightforward uphill battle for open-source LLMs. The landscape is constantly evolving: closed-source LLMs are updated by retraining on newer data regularly, open-source LLMs are released to catch up, and there is a myriad of evaluation datasets and benchmarks being used to compare LLMs, making singling out a best LLM especially challenging.
In this survey, we aim to consolidate recent studies on open-source LLMs and provide an overview of open-source LLMs that match or surpass ChatGPT in various domains. Our contributions are three-fold:
Who can benefit from this survey?
@@
In the following, we start by introducing background concepts (Section 2), then provide an in-depth review of open-source LLMs that beat ChatGPT in various domains (Section 3), followed by a discussion on insights and issues of open-source LLMs (Section 4), finally we conclude with a summary (Section 5).
{forest} for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, text width=2.7cm, s sep=6pt, calign=edge midpoint, , [LLM’s Capabilities, fill=gray!45, parent [General §3.1, domain_box [AlpacaEval, MT-bench, ELO rating, Open-LLM leaderboard, datasets [Llama-2, WizardLM, Zephyr, Deepseek, Yi, Mixtral, models] ] ] [Agent §3.2, domain_box [Using Tools, domain_box [API-Bank, ToolBench, APIBench, ToolAlpaca, etc., datasets [Gorilla, ToolLLaMA, few_models] ] ] [Self-Debugging, domain_box [InterCode-Bash, InterCode-SQL, MINT-MBPP, MINT-HumanEval, RoboCodeGen, etc., datasets ] ] [Following NL Feedback, domain_box [MINT, few_datasets] ] [Exploring Environments, domain_box [ALFWorld, InterCode-CTF, WebArena, datasets] ] [Lemur-chat, AgentLlama, OpenChat-3.5, models] ] [Logical Reasoning §3.3, domain_box [Math, domain_box [GSM8K, MATH, TheoremQA, etc., datasets [WizardMath, few_models] ] ] [Coding, domain_box [HumanEval, MBPP, APPs, etc., datasets [WizardCoder, few_models] ] ] ] [Modelling Long-context §3.4, domain_box [SCROLLS, Zero-SCROLLS, LongBench, L-Eval, BAMBOO, M4LE, etc., datasets [Llama-2-long, few_models] ] ] [Application-specific §3.5.5, domain_box [Query-focused Summarization, domain_box [QMSum, SQuALITY, CovidET, NEWTS, datasets [finetuned model, few_models] ] ] [Open-ended QA, domain_box [NQ,TriviaQA,NewsQA,SQuAD,Quoref, NarrativeQA,DROP, datasets [InstructRetro, few_models] ] ] [Medical, domain_box [Dreaddit, loneliness, MIMIC-CXR, OpenI, etc., datasets [MentaLLaMA, Radiology-Llama-2, few_models] ] ] [Generating Structured Data, domain_box [Rotowire, Struc-Bench-Latex, Struc-Bench-HTML, datasets [Struct-Bench, few_models] ] ] [Generating Critiques, domain_box [AlpacaFarm, FairEval, CritiqueEval, etc., datasets [Shepherd, few_models] ] ] ] [Trustworthiness §3.6, domain_box [Hallucination, domain_box [TruthfulQA, FactualityPrompt, FActScore, KoLA-KC, HaluEval, FACTOR, datasets [Platypus, Chain-of-Verification, etc., few_models] ] ] [Safety, domain_box [SafetyBench, XSTEST, etc., datasets] ] ] ]
@@ 2 Background
In this section, we briefly describe the fundamental concepts that relate to LLMs.
@@ 2.1 Training Regimes
All LLMs rely on large-scale self-supervised pre-training on Internet text data (Radford et al., 2018; Brown et al., 2020). Decoder-only LLMs follow the causal language modeling objective, through which the model learns to predict the next token conditioning on the sequence of previous tokens (Bengio et al., 2000). As per pre-training details shared by open-source LLMs (Touvron et al., 2023a), sources of text data include CommonCrawl [^5] https://commoncrawl.org, C4 (Raffel et al., 2020), GitHub, Wikipedia, books, and online discussion exchanges such as Reddit or StackOverFlow. It is widely acknowledged that scaling the size of pre-training corpus improves the model performance, and works hand-in-hand with scaling the model size, a phenomenon referred to as scaling laws, and analyzed in depth in (Hoffmann et al., 2022). Modern-day LLMs pre-train on a corpus from hundreds of billions to several trillions of tokens (Touvron et al., 2023b; Penedo et al., 2023).
@@
Fine-tuning aims to adapt a pre-trained LLM to downstream tasks, by updating weights with the available supervision, which usually forms a dataset orders of magnitude smaller than the one used for pre-training (Devlin et al., 2018). T5 (Raffel et al., 2020) was among the first to frame fine-tuning into a text-to-text unified framework, with natural language instructions describing each task. Instruction-tuning later extended fine-tuning by training jointly on several tasks (Wei et al., 2021; Aribandi et al., 2021), each described with natural language instructions. Instruction-tuning quickly gained in popularity, due to its ability to drastically improve zero-shot performance of LLMs, including on new tasks (unseen during training), and especially at larger models scale. Standard instruction-tuning with multi-task supervised fine-tuning (commonly known as SFT) may still not result in models that follow humans intentions while being safe, ethical and harmless, and can be futher improved with Reinforcement Learning from Human Feedback (RLHF): human annotators rank outputs from the fine-tuned model, which are used to fine-tune again with reinforcement learning (Ouyang et al., 2022b). Recent work showed that human feedback may be replaced with feedback from an LLM, a process referred to as Reinforcement Learning from AI Feedback (RLAIF) Bai et al. (2022b). Direct Preference Optimization (DPO) bypasses the need to fit a reward model to human preferences as in RLHF and instead directly fine-tunes the policy with a cross-entropy objective, achieving more efficient alignment of the LLM to human preferences.
A line of work focuses on quality over quantity when building an instruction-tuning dataset of diverse tasks: Lima (Zhou et al., 2023a) outperforms GPT-3 with a Llama-65B fine-tuned on just 1,000 examples, and Alpagasus (Chen et al., 2023c) improves on Alpaca (Taori et al., 2023) by cleaning its instruction fine-tuning dataset from 52k to 9k examples.
@@
Continual pre-training consists in performing another round of pre-training from a pre-trained LLM, typically with a lesser volume of data than in the first stage. Such process may be useful to quickly adapt to a new domain or elicit new properties in the LLM. For instance, continual pre-training is used in Lemur (Xu et al., 2023d) to improve coding and reasoning capacities, and Llama-2-long (Xiong et al., 2023) to extend context window.
@@
There exists several alternative methods for sequence generation with auto-regressive decoding with an LLM, which differ by the degree of randomness and diversity in the output. Increasing the temperature during sampling makes outputs more diverse, while setting it to 0 falls back to greedy decoding, which may be needed in scenarios necessitating deterministic outputs. Sampling methods top-k (Fan et al., 2018) and top-p (Holtzman et al., 2019) constrain the pool of tokens to sample from at each decoding step.
Several techniques aim to improve inference speed, especially at longer sequence length, which become problematic due to the attention complexity, which is quadratic with regards to input length. FlashAttention (Dao et al., 2022) optimizes reads/writes between levels of GPU memory, accelerating both training and inference. FlashDecoding (Dao et al., 2023) parallelizes the key-value (KV) cache loading in the attention mechanism, yielding a 8x end-to-end speedup. Speculative decoding (Leviathan et al., 2023; Chen et al., 2023b) uses an extra, small language model to approximate next token distribution from an LLM, which accelerates decoding without loss of performance. vLLM (Kwon et al., 2023) accelerates LLM inference and serving using PagedAttention, an algorithm for optimizing memory usage of attention keys and values.
@@ 2.2 Task Domains and Evaluation
3 Open-Source LLMs vs. ChatGPT
@@ 3.1 General Capabilities
As numerous LLMs are released week upon week, each claiming superior performance on certain tasks, it becomes increasingly challenging to identify true advancements and the leading models. Therefore, it is crucial to comprehensively assess the performance of these models across a broad spectrum of tasks to understand their general capabilities. This section covers benchmarks using LLM-based (e.g., GPT-4) evaluation and traditional (e.g., ROUGE (Lin, 2004) and BLEU (Papineni et al., 2002)) evaluation metrics.
  • MT-Bench (Zheng et al., 2023) is designed to test multi-turn conversation and instruction-following ability from eight perspectives: writing, roleplay, information extraction, reasoning, math, coding, knowledge I (STEM), and knoweldge II (humanities/social science). Stronger LLMs (e.g., GPT-4) are utilized as judges to evaluate the models for this benchmark.
  • AlpacaEval (Li et al., 2023e) is an LLM-based automatic evaluator based on AlpacaFarm (Dubois et al., 2023) evaluation set, which tests the ability of models to follow general user instructions. It benchmarks candidate models against Davinci-003 responses utilizing stronger LLMs (e.g., GPT-4 and Claude), which generate the candidate model’s win rate.
@@
  • FairEval-Vicuna (Wang et al., 2023d) utilizes both multiple evidence calibration and balanced position calibration on a set of 80 questions from the Vicuna Benchmark (Zheng et al., 2023). FairEval-Vicuna offers a more impartial evaluation outcome within the paradigm of adopting LLMs as evaluators, which closely aligns with human judgements.
Models MT-Bench AlpacaEval-2.0 Open LLM Leaderboard
Llama-2-70B-chat 6.86 13.87 -
WizardLM-70B 7.71 12.03 57.17
Zephyr-7B 7.34 10.99 52.15
Yi-34B - 29.66 69.42
Mixtral-8x7B 8.30 18.26 68.42
GPT-3.5-turbo 8.39 14.13 70.21
GPT-4 8.99 23.58 85.36
@@
Performance of LLMs
Llama-2-70B (Touvron et al., 2023b), a prominent open-source LLM from Meta, has been pre-trained on a massive dataset of two trillion tokens. It demonstrates remarkable results across various general benchmarks. In particular, Llama-2-chat-70B achieves a 13.87% win rate in AlpacaEval-2.0, approaching the performance of GPT-3.5-turbo . Zephyr-7B (Tunstall et al., 2023), a much smaller model, uses distilled direct preference optimization (Rafailov et al., 2023a) and achieves better results to 70B Llama-2 on MT-bench, scoring 7.34 against 6.86. Additionally, WizardLM-70B (Xu et al., 2023a) has been instruction fine-tuned using large amounts of instruction data with varying levels of complexity. It achieves a good score of 7.71 on MT-Bench. Although Zephyr-7B shows top performance in the MT-Bench, it falls short in the Open LLM Leaderboard, scoring only 52.15%.
Furthermore, Yi-34B pre-trained from scratch by developers at 01.AI  [^6] https://www.01.ai, stands out among all open-source LLMs with a remarkable score of 69.42% on . This performance is comparable to that of GPT-3.5-turbo, which scores 70.21%. Moreover, Yi-34B achieves a stunning score of 29.66% on AlpacaEval-2.0, surpassing GPT-4 with a significant margin. Mixtral(Jiang et al., 2024), a mixture of experts language model, is another top-performing open-source pretrained LLM, approaching GPT-3.5-turbo  performance on MT-Bench and Open LLM Leaderboard while outperforming GPT-3.5-turbo  on AlpacaEval-2.0.
Besides, UltraLlama (Ding et al., 2023) utilizes fine-tuning data with enhanced diversity and quality. It matches GPT-3.5-turbo’s performance in its proposed benchmark while exceeding it in areas of world and professional knowledge. Deepseek-67B(DeepSeek-AI et al., 2024), a recently released pretrained LLM with 15 more layers than Llama-2-70B, has demonstrated to surpass GPT-3.5-turbo on AlignBench(Liu et al., 2023a) – a comprehensive LLM benchmark in Chinese.